CORRESPONDENCES BETWEEN K3 SURFACES

Abstract. In this talk we show that there is a correspondence between some K3 surfaces with non-isometric transcendental lattices constructed as a twist of the transcendental lattice of the Jacobian of a generic genus 2 curve. Moreover, we show the existence of a correspondence between a general K3 surface with ρ = 17 and a Kummer surface having transcendental lattices Q-Hodge isomorphic. This reports a joint work with Giuseppe Lombardo and a geometric example due to Igor Dolgachev.